OpenStudy (anonymous):

please help...... ((2z-8)/(z^2-4))/((z^2+6z+8)/(z-4)) simplify completly..... show work

OpenStudy (anonymous):

do you know how to do this?

OpenStudy (dumbcow):

First Flip bottom fraction and Multiply Then factor and things will cancel \[2z-8 = 2(z-4)\] \[z^{2}-4 = (z+2)(z-2)\] \[z^{2}+6z+8=(z+4)(z+2)\]

OpenStudy (anonymous):

wow i forgot that the things cancelled.... okay what did you get for your answer?

OpenStudy (anonymous):

im not sure how you did that can you show me?

OpenStudy (dumbcow):

\[=\frac{2(z-4)(z-4)}{(z+2)(z+2)(z-2)(z+4)}\] hmm maybe they dont cancel in this case..haha

OpenStudy (anonymous):

yea and i think you factored wrong.....

OpenStudy (anonymous):

wait how did you factor that?

OpenStudy (dumbcow):

no everything is factored correctly

OpenStudy (dumbcow):

which one

OpenStudy (anonymous):

how did you do it then? i got something completly different but i thrust what you got....

OpenStudy (dumbcow):

are you sure its division and not multiplying the 2 fractions ??

OpenStudy (anonymous):

you flip the second one and change to multi.

OpenStudy (dumbcow):

correct

OpenStudy (anonymous):

and then can you show me what you did im completly lost

OpenStudy (dumbcow):

oh for the factoring part. well for 2z -8 i just pulled out a 2 from each term 2(z-4) = 2z - 8

OpenStudy (anonymous):

ok

OpenStudy (dumbcow):

For z^2 - 4, find factors of -4 that add up to 0 since there is no "z" term -2*2 = -4 and -2+2=0 (z-2)(z+2)

OpenStudy (anonymous):

wow thats sooo much easier now thanks :) can you help me with a couple more?

OpenStudy (dumbcow):

z^2 +6z +8 same thing, look for factors of 8 that add up to 6 4*2 =8 and 4+2 =6 (z+4)(z+2)

OpenStudy (dumbcow):

ok

OpenStudy (anonymous):

okay....... ((4)/(2x+1))-((3)/(2x))=

OpenStudy (anonymous):

can you show me step by step please?

OpenStudy (dumbcow):

adding/subtracting fractions you need to get common denominator 2x+1 cannot be factored 2x cannot be factored common denominator is 2x(2x+1) just like adding 1/2 + 1/3, common denominator is 2*3=6

OpenStudy (dumbcow):

now change numerators \[\frac{4}{2x+1} = \frac{(2x)(4)}{2x(2x+1)}\] \[\frac{3}{2x} = \frac{3(2x+1)}{2x(2x+1)}\] combine into 1 fraction \[\frac{(2x)(4) - 3(2x+1)}{2x(2x+1)}\]

OpenStudy (anonymous):

ok

OpenStudy (dumbcow):

multiply and add like terms on top \[\frac{8x -6x -3}{2x(2x+1)} = \frac{2x-3}{2x(2x+1)}\]

OpenStudy (anonymous):

can you help me with this one? ((m+5)/(2m^2-2))+((3)/(1-m))+((5)/(2m+2))

OpenStudy (dumbcow):

factor all the denominators 2m^2 -2 = 2(m-1)(m+1) 1-m = -(m-1) 2m+2 = 2(m+1) common denominator will include everything w/out repeats -2(m-1)(m+1)

OpenStudy (anonymous):

okay i can do it from there can you help me with this one? ((n+1)-(2/n))/((n+4)+(4/n))

OpenStudy (dumbcow):

\[=\frac{-(m+5)+(2(m+1))(3) + (-(m-1))(5)}{-2(m+1)(m-1)}\]

OpenStudy (anonymous):

can you help me with this one? please???? ((n+1)-(2/n))/((n+4)+(4/n))

OpenStudy (dumbcow):

ok combine fractions on top and bottom, then flip and multiply

OpenStudy (dumbcow):

common denominator will just be n

OpenStudy (anonymous):

ok

OpenStudy (dumbcow):

answer should be (n-1)/(n+2)

OpenStudy (anonymous):

yup that what i got are you good with summation notation and geometric means/sequences?

OpenStudy (dumbcow):

sure

OpenStudy (anonymous):

okay.... wellthe question is... insert four geometirc means between -7 and -224

OpenStudy (dumbcow):

hmm dont understand?

OpenStudy (anonymous):

nevermind i figured that one out......

OpenStudy (anonymous):

what about....

OpenStudy (anonymous):

\[\sum_{n=1}^{50}(1/4)(n+2)\]

OpenStudy (dumbcow):

\[=\frac{1}{4}\sum_{n=1}^{50}n+2 = \frac{1}{4}(\sum_{n=1}^{50}n +\sum_{n=1}^{50}2)\]

OpenStudy (anonymous):

im stuck on this one.. which term of the geometric sequence 243, -81, 27, . . . is (-1/9)?

OpenStudy (dumbcow):

8 th term

OpenStudy (anonymous):

how did you figure that out?

OpenStudy (dumbcow):

haha tried proving it using geometric sequence formula but got stuck when taking log anyway just continue the sequence of dividing by 3 and flipping the sign

OpenStudy (anonymous):

o so you flip the sign every other time? could you help me with this one? there are seven houses; in each are seven cats. each cat kills seven mice. each mouse would hvae eaten seven ears of wheat. each ear of wheat prodece seven measures of grain. how much grain is saved? i got 16807 grains saved

OpenStudy (dumbcow):

correct 7^3 mice 7^4 ears 7^5 grain

OpenStudy (anonymous):

ok that equals 16807 which is what i got:)

OpenStudy (anonymous):

okay change the repeating decimal .25 (sopposto have the bar above the 25) to an equivlant common fraction

OpenStudy (dumbcow):

?? why do you need to know that some fraction close to proportional of 1/4

OpenStudy (anonymous):

i dontk know its on my homework though ydo you know what it'd be?

OpenStudy (dumbcow):

haha i just guessed and put it in my calculator 25/99

OpenStudy (anonymous):

wow your good.... do you know how i would show that? just say guess and check? we turn this in for a grade thats why im asking

OpenStudy (dumbcow):

actually any n/99 = .nnnnnn

OpenStudy (anonymous):

in a certain credit union, money left on deposit for one year earns 4% intrest at the end ot the year. if you invested $100 at the beginning of each year in this credit union and did NOT withdraw the intrest due at the end of the year, how much would you hvae on deposit at the end of the tenth year?

OpenStudy (dumbcow):

=100(1.04 + 1.04^2+... +1.04^10) need sum of geometric sequence sum = a(1-r^n)/(1-r) sum = 1.04(1-1.04^10)/(1-1.04)

OpenStudy (dumbcow):

i get 1248.64

OpenStudy (anonymous):

so when you get your answer are you sopposto multiply it by 100?

OpenStudy (dumbcow):

yes

OpenStudy (anonymous):

okay just stay right there i have a couple more i just have to go and get another pencil mine just broke

OpenStudy (anonymous):

ok im back i dont know what formula you used

OpenStudy (dumbcow):

for finding sum of geometric sequence

OpenStudy (anonymous):

yea but can you tell me it with all of the variable in it please? i need to look and see if i have that one

OpenStudy (dumbcow):

\[s _{n} = \frac{a _{1}(1-r ^{n})}{1-r}\]

OpenStudy (anonymous):

i only have SofN=(Asub1-Asub1R^n)/(1-R) and... SofN=(Asub1-AsubnR)/(1-R) sooo.. which one should i use, and what numbers?

OpenStudy (dumbcow):

ues the first one, only difference is mine has factored out the a1 on top

OpenStudy (dumbcow):

a1 = first term r = common ratio

OpenStudy (anonymous):

ummmm... i should get the same answer right? i think im doing someting wrong im getting 1200.610712

OpenStudy (dumbcow):

1-(1.04)^10 = -.4802 1.04(-.4802) = -.49945 -.49945/(1-1.04) = 12.48635

OpenStudy (anonymous):

wait shouldnt it be..... (100-100(1.04)^10)/(1-1.04)

OpenStudy (dumbcow):

no im leaving the 100 on outside of sequence then we multiply the sum by 100

OpenStudy (anonymous):

but when i do it shouldnt i be getting the same thing? and i dont understand how your doing it

OpenStudy (dumbcow):

ok if you include the 100 then sequence will look like this =100(1.04) + 100(1.04^2) +...+100(1.04^10) a1 = 100(1.04) r = 1.04 sn = 100(1.04)(1-1.04^10)/(1-1.04)

OpenStudy (anonymous):

but my equation is Sn=(A1-A1R^n)/(1-R)

OpenStudy (dumbcow):

same thing factor out the A1

OpenStudy (anonymous):

can you do it my way when you factor out the A1 it confuses me

OpenStudy (anonymous):

....^because when you factor....

OpenStudy (dumbcow):

ok Sn = 100(1.04) -100(1.04)(1.04^10) / (1-1.04)

OpenStudy (anonymous):

ok ... a ball which rolls off a penthouse terrace falls 16 feet the first second, 48 feet the next second, and 80 feet the third second. if it continues to fall in this mannor, how far does it fall in the seventh second?

OpenStudy (dumbcow):

i believe this models a parabola of y=-16x^2 the change in y from 6 to 7 is how far it falls in 7th second 16(7^2) - 16(6^2) = 16(49-36) = 16(13) = 208

OpenStudy (anonymous):

a rubber ball dropped 40 feet rebounds on each bounce 2/5 of the distance from which it fell. how far will it travel before comming to a rest?

OpenStudy (dumbcow):

geometric sequence 40+40(2/5)+ 40(2/5)^2 +... looks like an infinite sequence before you get 0 lim n->inf (2/5)^n = 0 so in formula r^n = 0 a1=40 r=2/5 sum = 40 - 40(0) / 1-(2/5) sum = 40 /3/5 = 200/3 = 66.666

OpenStudy (anonymous):

what formula is that?

OpenStudy (dumbcow):

same as before sum of geometric sequence

OpenStudy (anonymous):

no but which one of mine because yours confuse me

OpenStudy (dumbcow):

\[s _{n}=\frac{a _{1}-a _{1}r^{n}}{1-r}\]

OpenStudy (anonymous):

what would i put for n?

OpenStudy (dumbcow):

infinity

OpenStudy (anonymous):

hu?

OpenStudy (dumbcow):

just substitute r^n=0

OpenStudy (anonymous):

would my answer be iin feet?

OpenStudy (dumbcow):

yes

OpenStudy (anonymous):

what does the graph for ..... look like? -cubic with one real solution and two complex solutions look like? -cubic with no real solutions? -quartic with no real solutions? -a quadratic with one real solution and one complex solution? -a quartic with two real solutions and two complex solutions? -a quadratic with no real solutions? -a quartic with no real solutions? -a cubic with 3 real solutions, but one is a double root? -a quartic with for real roots, but both are double roots?

OpenStudy (dumbcow):

umm number of real solutions represents number of x_intercepts cubic with no real solutions does not exist i believe